ImSonique:

Consider the function represented by the equation x – y = 3. What is the equation written in function notation, with x as the independent variable? f(x) = y + 3 f(x) = –y – 3 f(x) = –x + 3 f(x) = x – 3

jhonyy9:

so like a first step do you can separate the y in a part of this equation ? - better in the left part

ImSonique:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so like a first step do you can separate the y in a part of this equation ? - better in the left part \(\color{#0cbb34}{\text{End of Quote}}\) yes

jhonyy9:

so this is your equation x -y = 3 to separate the y on the left part like a first step you need subtract from both sides an x so in this way what will get ?

jhonyy9:

please collaborate

ImSonique:

f(x) = –y – 3?

jhonyy9:

no i ve said above subtract an x from both sides

ImSonique:

oh hehe sorry

ImSonique:

than would it be f(x) = y + 3

jhonyy9:

no so in this way x-x -y = 3 - x make the calcule pls

jhonyy9:

x-x = 0 -y = 3 - x but you need have on the left side +y so in this way you need multiplie both sides by -1 and will get (-y)(-1) = (-1)(3 - x) y = -3 +x or correct wrote y = x - 3 so f(x) = x - 3 hope helped understandably easy